OpenStudy (anonymous):

how would I write an equation of a line passing through the points (-2, -1) and (0, 4). Write the final answer in slope-intercept form.

OpenStudy (anonymous):

Well first find the slope, using the following formula: \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] where your coordinates represent \[(x_1,y_1)~~\text{and}~~(x_2,y_2)\] Once you manage to get that, plug in m and your first coordinates in the point - slope form: \[y-y _{1}=m(x-x _{1})\] do some little algebra and eventually you will have slope - intercept form \[y=mx+b\]

OpenStudy (anonymous):

K thank you very much!

OpenStudy (anonymous):

Y=2 1/2+2 does that answer look right?

OpenStudy (astrophysics):

Mhm, not quite

OpenStudy (astrophysics):

What is your slope?

OpenStudy (anonymous):

is it Y=2 1/2+5

OpenStudy (astrophysics):

No, what happened to your x?

OpenStudy (astrophysics):

Use iambatmans steps

OpenStudy (anonymous):

haha I did

OpenStudy (astrophysics):

Well lets do it together, step by step :), can you show me what you got for the slope?

OpenStudy (anonymous):

2 1/2

OpenStudy (astrophysics):

Right, but lets stick to 5/2

OpenStudy (astrophysics):

ok so now lets use point slope form, can you show me how you would do it

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (astrophysics):

Yeah, so plug in y1, x1 and m into the equation, leave y and x alone

OpenStudy (astrophysics):

\[y+1=5/2 (x+2)\] right? And we want y = mx+b

OpenStudy (anonymous):

y-(-1)=2 ½(x+2)

OpenStudy (astrophysics):

Yeah, that's fine, now we just do some algebra :)

OpenStudy (astrophysics):

so distribute 5/2 over (x+2) and then move +1 to the right! And we're done

OpenStudy (anonymous):

Y=5/2x+6

OpenStudy (astrophysics):

Subtract -1 not add :P

OpenStudy (astrophysics):

But good job!

OpenStudy (anonymous):

K thanks a lot for the help!

OpenStudy (misty1212):

HI!!

OpenStudy (astrophysics):

No problem :)

OpenStudy (misty1212):

did you get the slope?

OpenStudy (misty1212):

you already have the y intercept

OpenStudy (misty1212):

the points (-2, -1) and (0, 4)... since the point \((0,4)\) is on the graph, the y intercept is \(4\)

OpenStudy (astrophysics):

Yes, that's right, we could always use that to check our answer! :)

OpenStudy (misty1212):

?

OpenStudy (misty1212):

no checking involved find \(m\) and write \[y=mx+4\]

OpenStudy (astrophysics):

Yes, I know but it's good to use other methods and seeing what's going on as well.

OpenStudy (anonymous):

haha sorry but you guys' are confusing me:]

OpenStudy (astrophysics):

Don't worry about it haha, you can use any method, I usually prefer using what misty said after you understand the above method :P

OpenStudy (astrophysics):

As you get used to point slope form and slope intercept and see a relationship between them, it's pretty neat actually.

OpenStudy (anonymous):

K thanks!

OpenStudy (astrophysics):

Np, thank you Misty :)